# Semantic Operators: A Declarative Model for Rich, AI-based Analytics Over Text Data

### v2

@@ Abstract.
The semantic capabilities of language models (LMs) have the potential to enable rich analytics and reasoning over vast knowledge corpora. Unfortunately, existing systems lack high-level abstractions to perform bulk semantic queries across large corpora. We introduce semantic operators, a declarative programming interface that extends the relational model with composable AI-based operations for bulk semantic queries (e.g., filtering, sorting, joining or aggregating records using natural language criteria). Each operator can be implemented and optimized in multiple ways, opening a rich space for execution plans similar to relational operators. We implement our operators in LOTUS, an open source query engine with a DataFrame API. Furthermore, we develop several novel optimizations that take advantage of the declarative nature of semantic operators to accelerate semantic filtering, clustering and join operators by up to while offering statistical accuracy guarantees. We demonstrate LOTUS’ effectiveness on real AI applications including fact-checking, extreme multi-label classification, and search. We show that the semantic operator model is expressive, capturing state-of-the-art AI pipelines in a few operator calls, and making it easy to express new pipelines that achieve up to higher quality. Overall, LOTUS queries match or exceed the accuracy of state-of-the-art AI pipelines for each task while running up to 28 faster. LOTUS is publicly available at https://github.com/stanford-futuredata/lotus.
@@ 1. Introduction
The powerful semantic capabilities of modern language models (LMs) create exciting opportunities for building AI-based analytics systems that reason over vast knowledge corpora. Many applications require complex reasoning over large amounts of data, including both unstructured and structured data. For example a researcher reviewing recent ArXiv (arx, [n. d.]) preprints may want to quickly obtain a summary of relevant papers from the past week, or find the papers that report the best performance for a particular task and dataset. Similarly, a medical professional may automatically extract biomedical characteristics and candidate diagnoses from many patient reports (D’Oosterlinck et al., 2023). Likewise, organizations wish to automatically digest lengthy transcripts from internal meeting transcripts and chat histories to validate hypotheses about their business (dis, [n. d.]).
@@
Unfortunately, existing systems are insufficient for applications that require bulk semantic processing. Several systems provide logically row-wise LM operators, either in the form of a general-prupose LLM user-defined function (UDF) (ver, [n. d.]; dat, [n. d.]; noa, 2023; Liu et al., 2024b) or specialized row-wise operators (Liu et al., 2024c, a; Lin et al., 2024) for specific applications, such as conversational chat or data cleaning and extract-transform-load (ETL) tasks. These programming models fail to support a range of LM-based transformations across rows, such as ranking, grouping or joining records. Alternatively, many LM programming frameworks (lan, [n. d.]; lla, [n. d.]; Khattab et al., 2023) and research works (Khattab et al., 2021; Lee et al., 2019; Anantha et al., 2021; Gao et al., 2023; Izacard et al., 2022; Zelikman et al., 2022; Zhang et al., 2022) support combining LLM calls with retrieval-augmented generation (RAG) and Text2SQL, however these models are limited to point lookups to the data. RAG and Text2SQL methods first retrieve data using semantic search or LM-generated SQL calls, and then process the results with an LLM. Both are limited to analyses where an LLM can be fed a small number of documents or rows that are placed into a single model call’s context window (e.g., questions answered in one document in the corpus). However, bulk semantic processing applications may involve analysis across multiple rows, for clustering, ranking or joining an entire dataset.
Towards a declarative programming interface for bulk semantic processing, we propose semantic operators, which extend the relational model with AI-based operations that users can compose into powerful, reasoning-based query pipelines over structured and unstructured data. These operators include semantic filters, joins, rankings, aggregations, and projections, which take natural language expressions given by the programmer. To provide an implementation of these operators, we present the LOTUS system, which exposes these operators through a simple DataFrame-based programming interface. LOTUS’ query engine efficiently and accurately executes queries with semantic operators, while abstracting away low-level details like model context length limits and algorithmic choices. Finally, we propose novel optimizations for the semantic filter, join, ranking and groupby operators that can improve performance by up to over a naive, expensive implementation while guaranteeing accuracy similar to the naive implementation.
| Operator | Description |
| Returns the tuples in a table that pass the provided langex predicate. | |
@@
| Joins a table against a second table by keeping all pairs of tuples that pass the provided langex predicate. | |
| Performs an aggregation over the input tuples according to the langex, which specifies a commutative, associative aggregation function over a list of tuples. | |
| Ranks each tuple and returns the best according to the langex, which specifies a ranking function that sorts a list of tuples. | |
@@
| Groups the tuples into categories based on the langex, which specifies a grouping criteria. | |
| Performs a projection, returning a new column, according to the provided langex. | |
@@
Figure 1 begins to demonstrate the power of LOTUS’ declarative programming model and optimized query engine. For a fact-checking task on the FEVER dataset (Thorne et al., 2018), we write an intuitive LOTUS program in less than 50 lines of code by composing semantic operators (i.e., filters, maps, and joins). The modularity and composability of these semantic operators make it easy to quickly write programs and explore the design space on this task. In doing so, we find the un-optimized LOTUS programs can reproduce and improve accuracy on this task by percentage points compared to a recent state-of-the-art fact-checking pipeline, FacTool (Chern et al., 2023), which was written in over 750 lines of code. LOTUS also provides query efficiency by default, yielding lower execution time than the FacTool implementation.
LOTUS’ optimizer exploits the rich implementation design space of semantic operators to provide novel optimizations for bulk semantic operations. To demonstrate the power of programming with semantic operators, we implement optimizations for several costly operators, i.e., semantic filter, join, ranking and group-by. Specifically, we develop a novel, lossless optimization for ranking. Additionally, for filters, joins and group-bys, we define and implement both a tractable, but expensive, reference algorithm as well as an approximation algorithms with probabilistic guarantees on accuracy with respect to the reference algorithm. Our approximation algorithms leverage build on statistical techniques used in prior works (Kang et al., 2021, 2020) with novel and efficient proxy scorers that leverage small LMs or semantic embeddings. Figure 1 demonstrates the effectiveness of these methods. The program implemented using LOTUS’ optimizations with recall and precision targets and a error probability attains high accuracy, of the unoptimized program’s accuracy, and runs faster than the unoptimized program.
We systematically evaluate the expressiveness of the semantic operator model and the efficiency of our algorithms and optimizations with statistical accuracy guarantees through several real applications, including fact-checking, extreme multi-label classification, and search, which had previously been implemented by hand in AI research papers. We show that LOTUS’ programming model is highly expressive, capturing high quality and state-of-the-art query pipelines with low development overhead for these wide-ranging applications. Furthermore, our algorithms and optimizations can speed up query execution by up to while offering statistical guarantees on accuracy.
Specifically, on the FEVER dataset (Thorne et al., 2018) for fact-checking, LOTUS programs can reproduce a recent state-of-the-art pipeline (Chern et al., 2023), as shown in Figure 1, in few lines of code, and implement a new pipeline with a simple change of operators that improves accuracy by , while offering or lower execution time with or without batching, respectively, while providing statistical accuracy guarantees. In the extreme multi-label classification task on the BioDEX dataset (D’Oosterlinck et al., 2023), LOTUS reproduces state-of-the art result quality (D’Oosterlinck et al., 2024) with it’s join operator, while providing an efficient algorithm that provides lower execution time than the naive algorithm, while again offering statistical guarantees on accuracy. In the search and ranking application, LOTUS allows a simple composition of operators to achieve higher nDCG@10 than the vanilla retriever and re-ranker, while also providing query efficiency, with lower execution time than LM-based ranking methods (Qin et al., 2024) used by prior works.
Our main contributions are the following:
We propose semantic operators as a programming model for AI-based analytics. Semantic operators (e.g., filters, joins, top-k ranking, aggregations, and group-bys) are declarative and can be implemented in multiple ways, creating a rich design space for execution plans.
We present efficient algorithms and query optimizations for several costly operators, i.e., semantic filter, join, top-k ranking and group-by. We develop a novel, lossless optimization for ranking and approximations for filters, joins and group-bys that provide statistical accuracy guarantees with respect to tractable, but expensive, reference algorithms.
We provide a reference implementation of semantic operators and optimizations in the open-source LOTUS system.
2. Semantic Operators
We define semantic operators as declarative transformations on one or more relations, parameterized by a natural language expression, that can be implemented by a variety of AI-based algorithms. Semantic operators provide a seamless extension to the relational model, operating over tables that may contain traditional structured data as well as unstructured fields such as free-form text. These new operators can be composed with traditional relational operators to build rich queries that can be transparently optimized, which we discuss in Section 4. Table 1 lists a core set of semantic operators we propose in this paper, which cover the common semantic transformations we observe in real-world applications and mirror the key transformation patterns in standard relational operators. Specific analytics systems may of course provide additional semantic operators or APIs that combine the ones we discuss here.
@@ 2. Semantic Operators
Each semantic operator takes a parameterized natural language expressions (langex for short), which are natural language expressions that specify a function over one or more attributes for the given semantic transformation. The langex signature depends on its semantic operator and may represent a predicate, aggregation, comparator function, or projection in natural language.
Figure 2 shows an example of using semantic operators in LOTUS’ dataframe API to load data about ArXiv papers and perform a summarization task, which finds relevant papers using a top-k search, filters according to whether the paper claims to outperform a specific baseline, and then summarizes the remaining papers into a single digest with an aggregation. The code in Figure 2 highlights various langex signatures: the sem_filter langex provides a predicate that indicates a filter criteria to apply over the title and abstract attributes, while sem_agg takes a langex that provides an associative aggregation expression, which here indicates a many-to-one summarization task over abstracts.
Together, each langex and semantic operator specifies a modular semantic transformation, whose behavior is defined with respect to a world model , which may be a powerful LLM that captures a probability distribution over the vocabulary .
We briefly overview several semantic operators, both ones whose definitions directly extend their analogs in traditional relational algebra, and those with more complex definitions.
@@
Semantic Filter is a unary operator over the relation and returns the relation , where the natural language predicate is specified by the langex with respect to model . The langex is paramterized by one or more attributes from the relation that are relevant for evaluating the predicate.
@@
Semantic Join provides a binary operator over relations and to return the relation , once again taking the langex with respect to model . Here the langex is parameterized by the left and right join keys and describes a natural language predicate over both.
@@
Semantic Top-k imposes a ranking111This definition implies that imposes a total and consistent ordering. However, this definition can also be softened to assume partial orderings and noisy comparisons with respect to model . over the relation and returns . Here the signature of the langex provides a general ranking criteria according to one or more attributes of the input relation. The underlying system can use this langex to impose a ranking over any subset of rows, according to the chosen implementation.
@@
Semantic Aggregation performs an aggregation over the input relation, returning . Here, the the langex signature provides a commutative and associative aggregation function, which can be applied over any subset of rows to produce an intermediate results. We note that the langex with respect to model assumes infinite context and can take an arbitrary number of input tuples. Managing finite context limits of the underlying model is an implementation detail left to the system. Semantic aggregations can be useful for many tasks involving many-to-one reduction patterns, such as summarization or question-answering over multiple documents.
@@
Semantic Group-by takes a langex that specifies a grouping criteria and a target number of groups. As a example extending that of Figure 2, we might group-by the topics presented in each paper abstract with a target of 10 group. Here, the group-by operator must discover representative group labels and assign each tuple a label. In general, performing the unsupervised group discovery is a clustering task, which is NP-hard (Dasgupta, [n. d.]). Clustering algorithms over points in a metric space typically optimize the potential function tractably using coordinate descent algorithms, such as k-means. In the semantic group-by, the clustering task is over unstructured documents with a natural language similarity function specified by the user langex , which imposes a real-valued score between a tuple and a candidate label , and the operator poses the following optimization problem:

### v3

# Semantic Operators: A Declarative Model for Rich, AI-based Data Processing

@@ Abstract.
The semantic capabilities of large language models (LLMs) have the potential to enable rich analytics and reasoning over vast knowledge corpora. Unfortunately, existing systems either empirically optimize expensive LLM-powered operations with no performance guarantees, or serve a limited set of row-wise LLM operations, providing limited robustness, expressiveness and usability. We introduce semantic operators, the first formalism for declarative and general-purpose AI-based transformations based on natural language specifications (e.g., filtering, sorting, joining or aggregating records using natural language criteria). Each operator opens a rich space for execution plans, similar to relational operators. Our model specifies the expected behavior of each operator with a high-quality gold algorithm, and we develop an optimization framework that reduces cost, while providing accuracy guarantees with respect to a gold algorithm. Using this approach, we propose several novel optimizations to accelerate semantic filtering, joining, group-by and top-k operations by up to . We implement semantic operators in the LOTUS system and demonstrate LOTUS’ effectiveness on real, bulk-semantic processing applications, including fact-checking, biomedical multi-label classification, search, and topic analysis. We show that the semantic operator model is expressive, capturing state-of-the-art AI pipelines in a few operator calls, and making it easy to express new pipelines that match or exceed quality of recent LLM-based analytic systems by up to , while offering accuracy guarantees. Overall, LOTUS programs match or exceed the accuracy of state-of-the-art AI pipelines for each task while running up to faster than the highest-quality baselines. LOTUS is publicly available at https://github.com/lotus-data/lotus.
@@ 1. Introduction
The powerful semantic capabilities of modern language models (LLMs) create exciting opportunities for building AI-based analytics systems that reason over vast knowledge corpora. Many applications require complex reasoning over large amounts of data, including both unstructured and structured data. For example a researcher reviewing recent ArXiv (arx, [n.d.]) preprints may want to quickly obtain a summary of relevant papers from the past week, or find the papers that report the best performance for a particular task and dataset. Similarly, a medical professional may automatically extract biomedical characteristics and candidate diagnoses from many patient reports (D’Oosterlinck et al., 2023). Likewise, organizations wish to automatically digest lengthy transcripts from internal meeting transcripts and chat histories to validate hypotheses about their business (dis, [n.d.]).
@@
Unfortunately, existing systems are insufficient for bulk semantic processing, either limiting their expressiveness to row-wise LLM execution models or focusing on empirical optimizations without any formalism or accuracy guarantees, leading to unexpected behaviors and frequent failures. First, several systems provide logically row-wise LLM operators, either in the form of a general-purpose AI user-defined functions (UDF) (MotherDuck, [n.d.]; WilliamDAssafMSFT, 2024; noa, [n.d.]d; ver, [n.d.]; dat, [n.d.]; noa, 2023; Liu et al., 2024a) or specialized row-wise operators (Liu et al., 2024c, b; Lin et al., 2024) for specific applications, such as conversational chat, data cleaning and extract-transform-load (ETL) tasks. These programming models fail to support a range of LLM-based transformations across rows, such as ranking, grouping or joining records. Alternatively, more recent LLM-based analytics systems, such as DocETL (Shankar et al., 2024) and UQE (Dai et al., 2024), empirically optimize LLM-based operations with no accuracy guarantees, lacking any formalism to define correct behavior, which hinders their robustness and usability, as we show in Section 5. These obstacles highlight a core challenge of integrating semantic-based processing within reliable query systems due to the inherent ambiguity of natural language instructions and LLM outputs.
Towards robust, declarative query systems for bulk semantic processing, we propose the semantic operator model, which extends the relational model with AI-based operations. Semantic operators provide the first formalism for general-purpose and multi-row, AI-based transformation, including semantic filters, joins, top-k rankings, aggregations, and projections. Each operator takes a concise natural language signature, given by the programmer, and its behavior is fully specified by a tractable, high-quality ”gold” algorithm, which indicates how to orchestrate the underlying AI model over the data. Our optimization approach then exploits the rich design space of semantic operator execution plans to reduce cost, while providing statistical accuracy guarantees with respect to the gold algorithm (i.e., guarantees that the output of the optimized operator will be similar to that of the gold algorithm). We propose several novel optimizations for the semantic filter, join, top-k and group-by operators. Our methods include approximation algorithms, building on statistical techniques used in prior works (Kang et al., 2021, 2020) with novel and efficient proxy scorers using small LLMs or semantic embeddings. We implement semantic operators in LOTUS (LLMs Over Tables of Unstructured and Structured data), an open source system that exposes these operators through a simple DataFrame-based programming interface.
| Operator | Description | Definition | Gold Algorithm |
| Returns the tuples that pass the langex predicate. | Compute | ||
@@
| Joins a table against a second table by keeping all tuple pairs that pass the langex predicate. | Compute | ||
| Aggregates input tuples according to the langex reducer function. | Perform a reduce algorithm, recursively computing , | ||
| Returns an ordered list of the best tuples according to the langex ranking criteria. | Perform top-k sorting algorithm using pairwise comparisons, | ||
@@
| Groups the tuples into categories based on the langex grouping criteria. | Obtain centers with a clustering algorithm, and perform pointwise assignments | ||
| Performs the projection specified by the langex. | Compute | ||
@@
Figure 1 illustrates the power of LOTUS’ declarative programming model and optimized query engine. For a fact-checking task on the FEVER dataset (Thorne et al., 2018), we write an intuitive LOTUS program to re-implement a recent state-of-the-art fact-checking pipeline, FacTool (Chern et al., 2023), which was originally implemented in over 750 lines of code. Our un-optimized LOTUS program composes semantic operators (filters, maps and joins) and achieves higher accuracy on this task. LOTUS also provides query efficiency by default, yielding lower execution time than the original FacTool implementation. Furthermore, the optimized LOTUS program, configured with accuracy targets and a allowed error probability, runs faster and obtains accuracy compared to the un-optimized LOTUS program.
We systematically evaluate LOTUS on four real bulk-semantic processing applications: fact-checking, biomedical multi-label classification, search, and topic analysis. These include applications expressible and studied by recent LLM-based analytics systems (Shankar et al., 2024; Dai et al., 2024), where we compare against those systems, as well as others, where we compare against hand-designed pipelines from the AI literature. We demonstrate that the semantic operator model is effective, capturing state-of-the-art AI pipelines in a few operator calls that reliably match or exceed quality of recent LLM-based analytics systems by up to , while offering accuracy guarantees and running up to faster than the best-quality baselines. Furthermore, our gold algorithms provide high-quality results, and our optimizations speed up query execution by up to with respect to them. Specifically, our LOTUS program re-implements FacTool’s recent state-of-the-art fact-checking pipeline (Chern et al., 2023), as shown in Figure 1, and achieves higher accuracy with lower execution time on the FEVER dataset (Thorne et al., 2018), outperforming alternative LLM-powered analytics systems on both accuracy and efficiency. In the extreme multi-label classification task on the BioDEX dataset (D’Oosterlinck et al., 2023), LOTUS reproduces state-of-the art result quality (D’Oosterlinck et al., 2024) and the LLM-powered analytics baselines, while providing accuracy guarantees and speedups with respect to our gold algorithm. In the search and ranking application, LOTUS allows a simple composition of operators to achieve higher nDCG@10 than the next best performing baselines, while also providing query efficiency, with lower execution time than alternative high-quality algorithms. In the topic analysis task, LOTUS efficiently processes hundreds of ArXiv papers to discover a taxonomy of key topics in dozens of seconds, while once again providing configurable speedups with statistical accuracy guarantees.
Overall, our main contributions are the following:
We propose semantic operators, the first formalized programming model for general-purpose AI-based transformations with natural language specifications.
We define an optimization framework for semantic operators with statistical accuracy guarantees, and novel optimizations for several operators, including semantic filters, joins, top-k and group-by, yielding up to speedups.
We extensively evaluate the expressiveness of the semantic operator model, the quality of our gold algorithms, and the benefit of our optimizations, including comparisons to DocETL, UQE, and AI UDFs.
2. The Semantic Operator Model
Semantic operators provide a declarative interface for semantic-based manipulation and access to data. Extending the data independence model of relational systems (Codd, 1970), these new AI-based operators introduce model-data independence, which separates the application logic from the underlying ML-based algorithm and access pattern that specifies how data is passed to model invocations. Unlike relational operators, semantic operators are parametrized by natural language expressions and rely on AI-based computation, making their behavior inherently ambiguous and imprecise. We address this key challenge by presenting a formalism for defining semantic operators, their behavior and correct optimizations, in this section. In Section 3, we build on this formalism to provide optimized execution plans with accuracy guarantees, similar to relational query optimizers.
2.1. Defining Semantic Operators
Definition. A semantic operator is a declarative transformation over one or more datasets, parameterized by a natural language expression. Each semantic operator can be implemented by many AI-based algorithms, and its behavior is defined with respect to a gold algorithm.
Table 1 lists a core set of semantic operators, which cover common semantic transformations in real-world applications and mirror key transformations in relational operators.
Specific systems may, of course, provide additional semantic operators beyond the ones we discuss here.
@@ 2.1. Defining Semantic Operators
Each semantic operator takes a parameterized natural language expressions (langex for short), which are natural language expressions that specify a function over one or more attributes. As Figure 2 demonstrates, the langex signature varies for different semantic transformations. The figure shows a LOTUS program,
which performs a sem_join, sem_map and sem_agg over datasets of research papers and user projects.
While the sem_join langex signature provides a natural language predicate, the sem_map langex signature indicates a natural language projection, and the sem_agg langex is a commutative, associative aggregator expression, which here indicates a summarization task over abstracts.
We provide a high-level definition of each semantic transformation with respect to the user langex and a world model111In practice, the world model, may be the strongest LLM a practioner has available., , which captures a probability distribution over the vocabulary . For example, semantic filter returns , where is the input relation and represents a natural language predicate evaluated on tuple with model M. Notably, the definition of each semantic operator can be implemented by multiple AI-based algorithms, and different decisions as to how to invoke the model over the relation have consequences on the algorithm’s result quality. Thus, to fully specify the ground truth behavior of each semantic operator, we define a gold algorithm, which is a computable and tractable ML algorithm that produces high-quality results and specifies the model’s access pattern over the data, .
2.2. Defining Correct Optimizations for Semantic Operators
Semantic operators create a rich design space of diverse execution plans. While gold algorithms define a tractable, high-quality implementation, they are often expensive, with the complexity of LLM calls scaling linearly or quadratically in the dataset cardinality. For example, the semantic filter’s gold algorithm entails invoking the LLM in a linear pass over each tuple. However, an alternate AI-based algorithm may reduce cost, while offering close results. In the case of semantic filters, close results, can be measured according to accuracy metrics, in this case recall and precision, and guaranteed with high probability compared to the accuracy obtained by the gold algorithm. As such, we define correct optimizations for semantic operators222Our optimization approach can naturally extend to multi-operator execution plans by composition. In multi-operator programs, the user may configure an accuracy target and failure probability per operator, or for the full program, in which case the optimizer assigns a target accuracy and failure budget to each operator. below.
Definition. A correct optimization for a given semantic operator, and a gold algorithm for that operator, reduces cost while providing statistical accuracy guarantees with respect to the gold algorithm. Specifically, the optimization should ensure an accuracy target, , is met with probability .
In general, optimized semantic operator plans allows for both lossless optimizations and approximations with respect to a gold algorithm. This formulation builds on approaches in approximate query processing and assumes some error is often tolerable, which we find (Section 5) is reasonable for achieving high-quality results, for semantic processing, which is inherently non-exact.
2.3. Core Semantic Operators
We now overview several core semantic operators, providing the operator’s definition and at least one gold algorithm for each. We discuss design decisions for our gold algorithms based on state-of-the-art algorithms studied in the AI literature, well-known failure cases, such as long-context challenges (Liu et al., 2023), and our empirical evaluation in Section 5.
@@ 2.3. Core Semantic Operators
Semantic Filter is a unary operator over the relation and returns the relation , where the langex provides a natural language predicate over one or more attributes.
Gold Algorithm. Our gold algorithm runs batched LLM calls over all tuples in relation . Each model invocation, , prompts the LLM with a single tuple , the langex predicate, and an operator-specific instruction to generate a boolean value. This simple choice avoids well-studied long-context issues (Liu et al., 2023) by processing rows independently rather than batching multiple tuples within a single prompt invocation.
@@
Semantic Join provides a binary operator over relations and to return the relation . Here the langex is parameterized by the left and right join keys and describes a natural language predicate over both.
Gold Algorithm. The gold algorithm implements a nested-loop join pattern, performing a single predicate evaluation over a tuple pair with each model invocation , similar to filters This yields an LM call complexity, where and are the table sizes of the left and right join tables respectively. This quadratic join algorithm is suitable for small join tables but scales poorly for large tables.
@@
Semanic Top-k imposes a ranking333This definition implies that imposes a total and consistent ordering. However, this definition can also be softened to assume partial orderings and noisy comparisons with respect to model . over the relation and returns the ordered sequence, . Here, the langex signature provides a general ranking criteria according to one or more attributes.
Gold Algorithm. Two important algorithmic design decisions for the semantic top-k include how to implement LLM-based comparison and how to aggregate ranking information from these comparisons. Our gold algorithm uses pairwise LLM comparisons for the former, and a quick-select top-k algorithm (Hoare, 1961) for the latter. We briefly describe the reason for these choices and alternatives considered. Our decisions build on prior works, which have studied LLM-based passage re-ranking (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Ma et al., 2023; Pradeep et al., 2023a, b; Qin et al., 2024; Sachan et al., 2022; Sun et al., 2023) and ranking with noisy comparisons (Shah and Wainwright, 2016; Braverman and Mossel, [n.d.]) with the goal of achieving high quality results in a modest complexity of LLM calls or comparisons.
First, pairwise-prompting methods offer a simple and high-quality approach that feeds a single pair of tuples to each LLM invocation, , prompting the model to compare the two inputs and output a binary label. The two main classes of alternatives are point-wise ranking methods (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Sachan et al., 2022; Wu et al., 2024), and list-wise ranking methods (Ma et al., 2023; Pradeep et al., 2023a, b; Sun et al., 2023), both of which have been shown to face quality issues (Sun et al., 2023; Qin et al., 2024; Desai and Durrett, 2020). We verify these limitations in Section 5. In contrast, pairwise comparisons have been shown to be effective and relatively robust to input ordering (Qin et al., 2024).
In addition, we consider several possible rank-aggregation algorithms, including quadratic sorting algorithms, a heap-based top-k algorithm and a quick-select-based top-k ranking algorithm. All three alternatives represent candidate gold algorithms. Our evaluation in Section 5 demonstrates that each of these sorting algorithms yield high-quality results, comparable to one another. However, the quick-select-based algorithm offers an efficient implementation with at least an order of magnitude fewer LLM calls than the quadratic sorting algorithm and more opportunities for efficient batched inference, leading to lower execution time, compared to a heap-based implementation. The quick-select top-k algorithm proceeds in successive rounds, each time choosing a pivot, and comparing all other remaining tuples to the pivot tuple to determine the rank of the pivot. Because each round is fully parallelizable, we can efficiently batch these LLM-based comparisons before recursing in the next round.
@@
Semantic Aggregation performs a many-to-one reduce over the input relation, returning . Here, the the langex signature provides a commutative and associative aggregation function444We note that the ordering of inputs within an LLM prompt invocation can in fact affect results quality for some tasks. Thus, programmers may wish to override the commutativity and associativity assumptions of the semantic aggregation langex. For this, the LOTUS API exposes a partioner function which groups and orders inputs within LLM prompts according to an arbitrary user-specification., which can be applied over any subset of rows to produce an intermediate results. We note that the langex iteself is model-agnostic, assuming infinite context. Managing finite context limits of the underlying model is an implementation detail left to the system.
Gold Algorithm. Our gold algorithm must efficiently orchestrate the LLM to manage long context inputs, which may degrade result quality (Liu et al., 2023) or overflow the underlying model’s context length. We select a hierarchical reduce pattern for the gold algorithm. Our choice builds on the LLM-based summarization pattern studied by prior research works (Wu et al., 2021; Chang et al., 2024; Adams et al., 2023) and deployed systems (lla, [n.d.]; lan, [n.d.]), which we briefly overview.
Prior works primarily implement one of two aggregation patterns: either a fold pattern, which performs a sequential, linear pass over the data, while iteratively updating an accumulated partial answer, or a hierarchical reduce pattern, which recursively aggregates the input data to produce partial answers until a single answer remains. Both represent candidate gold algorithms, however, the hierarchical pattern has been shown to produce higher quality results for commutative, associative aggregation tasks, like summarization, in prior work (Chang et al., 2024) and allows for greater parallelism during query processing, making it our default choice.
@@
Semantic Group-by takes a langex that specifies a projection from a tuple to an unknown group label, as well as a target number of groups, which specifies the desired granularity of group labels. As a example, a user might group-by the topics presented in a set of ArXiv papers, wishing to find 10 key groups. The group-by operator must discover representative group labels and assign a label to each each tuple. In general, performing the unsupervised group discovery is a clustering task, which is NP-hard (Dasgupta, [n.d.]). Clustering algorithms over points in a metric space typically optimize the potential function tractably using coordinate descent algorithms, such as k-means. For the semantic group-by, the clustering task is over unstructured fields with a natural language similarity function specified by , which imposes a real-valued score between a tuple and a candidate label . This operator poses the following optimization problem: